# Forging Tokens for Improved Storage-efficient Training

### v1

@@ Abstract
Recent advancements in Deep Neural Network (DNN) models have significantly improved performance across computer vision tasks. However, achieving highly generalizable and high-performing vision models requires extensive datasets, leading to large storage requirements. This storage challenge poses a critical bottleneck for scaling up vision models. Motivated by the success of discrete representations, SeiT proposes to use Vector-Quantized (VQ) feature vectors (i.e., tokens) as network inputs for vision classification. However, applying traditional data augmentations to tokens faces challenges due to input domain shift. To address this issue, we introduce TokenAdapt and ColorAdapt, simple yet effective token-based augmentation strategies. TokenAdapt realigns token embedding space for compatibility with spatial augmentations, preserving the model’s efficiency without requiring fine-tuning. Additionally, ColorAdapt addresses color-based augmentations for tokens inspired by Adaptive Instance Normalization (AdaIN). We evaluate our approach across various scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, robustness benchmarks, and ADE-20k semantic segmentation. Experimental results demonstrate consistent performance improvement in diverse experiments, validating the effectiveness of our method. Code is available at https://github.com/naver-ai/tokenadapt.
@@ 1 Introduction
Significant advances in Deep Neural Network (DNN) models [16, 48] have substantially improved their performance across various computer vision tasks. However, achieving highly generalizable and high-performing vision models demands an extensive dataset, often reaching billions of data points [25, 44, 34] with large storage requirements (e.g., the LAION-5B [42] dataset requires a 240TB storage size). This storage challenge emerges as a critical bottleneck for scaling up vision models. Consequently, optimizing storage efficiency becomes important for the practical advancement of scaled-up vision models.
@@
Images require more storage than texts, due to their intrinsic complexity in converting continuous natural light signals into discrete pixels for computer representation. However, a considerable portion of image data corresponds to redundant or unnecessary details, such as repeated patterns or imperceptible details [39]. These properties of image data lead to large storage requirements and also induce negative impacts on visual recognition tasks as mentioned in [10, 21, 33, 1, 20, 43].
To build a deep representation away from redundant and superfluous information in image details, earlier researches [49, 17, 56, 3, 2] have explored ways to extract condensed information from images. One notable study, VQ-VAE [49], emphasized the importance of discrete representations, suggesting it as a potentially more natural fit for various modalities. They demonstrated that an image can be sufficiently expressed as a combination of vector-quantized (VQ) feature vectors by achieving similar performance as their continuous counterparts in the generation task. Recent studies [17, 39] further support this claim by successfully synthesizing high-resolution images from discrete feature vectors. Building on this understanding, SeiT [36] delved into optimizing storage efficiency by utilizing discrete representations (i.e., tokens) as network inputs. SeiT suggests using the vector-quantized feature instead of image pixels, which substantially reduces the storage size for a dataset by excluding redundancies in images.
These findings highlight the effectiveness of leveraging pre-extracted tokens. They successfully capture the essence of original pixel data while taking advantage of the inherent storage efficiency in the VQ technique. Nevertheless, this approach introduces a significant challenge: applying existing data augmentations designed for pixel images to token embedding is hindered by input domain shift. Previous studies [12, 11, 48] have highlighted the crucial role of data augmentation in model performance, positioning it as a fundamental building block extensively employed in various training recipes. Considering its significance in DNN-based models, addressing the challenges posed by the application of data augmentation to token-based vision models is crucial. This resolution holds the key to overcoming major performance bottlenecks, thereby enhancing the effectiveness of these models across diverse vision tasks, including semi-, self-supervised learning, and limited data scenarios.
To tackle this issue, we propose two simple yet effective augmentation strategies called TokenAdapt and ColorAdapt. We designed TokenAdapt motivated by the inherent property of token embedding that encapsulates image coordinate information and can be decoded back to the original image space. The purpose of our TokenAdapt is to realign the token embedding space for better compatibility with augmentation. We achieve this by converting token embedding into a learned feature space suitable for pixel-based augmentation techniques, applying the augmentations, and then reverting the augmented feature back to the original token embedding space. In this way, we can reuse established image-augmentation policies as-is. Besides, it allows us to keep the pre-trained weights of the original token-based model unchanged, avoiding significant computational overhead such as fine-tuning.
While TokenAdapt can handle geometric data augmentations, applying color-based data augmentations, which typically operate on pixel values, to a non-pixel domain remains a non-trivial task. Here, we introduce ColorAdapt, a method designed to enable controllable color changes for token embedding. Inspired by Adaptive Instance Normalization (AdaIN [24]), ColorAdapt aims to preserve object structure while introducing significant color variations. In the training phase, we employ TokenAdapt and ColorAdapt for data augmentation. With them, we find that our strategy can improve the effectiveness of token-based training in terms of model performance.
@@
We validate the effectiveness of our proposed token-based augmentation strategies through experiments across various scenarios where data augmentation is crucial: (1) storage-efficient ImageNet-1k classification, (2) transfer learning on fine-grained classification, (3) evaluation on robustness benchmarks, and (4) ADE-20k semantic segmentation. As depicted in Figure 1, our method consistently outperforms comparison methods in storage-efficient ImageNet-1k classification with limited storage sizes. These results confirm that our proposed method significantly contributes to the storage efficient model by augmenting data effectively. Furthermore, we showcase how our augmentation strategies enhance the generalizability of trained models, as evidenced by our experimental results on fine-grained classification and robustness evaluation. Lastly, our method’s applicability to an alternative input format (e.g., DCT coefficients) demonstrates its extensibility.
@@ 2 Related Work
Learning with tokenization. The use of discrete tokens for training vision models has recently gained attention, particularly in self-supervised representation learning. BeiT [4] employs the Masked Image Modeling (MIM) framework to recover discrete tokens from masked images and CIM [18] utilizes tokens in Corrupted Image Modeling. MAGE [30] demonstrates the incorporation of tokens from VQGAN [17] for both generative training and representation learning within a unified training framework. However, existing studies mainly focus on self-supervised representation learning and rely on online tokenization, converting images to tokens during each iteration of model training. This approach incurs notable memory consumption and computational costs, limiting the full exploitation of the storage efficiency of discrete tokens.
@@
Storage-efficient vision training. Storage-efficient vision training aims to reduce the dataset storage size. To this end, dataset distillation (DD) [53] creates a condensed dataset by leveraging the knowledge of the original dataset. Recent research [29, 40, 41, 59, 60] highlights the effectiveness of dataset distillation in improving training efficiency. However, its high computational complexity poses challenges, particularly when applied to large-scale datasets like ImageNet-1k [14]. Sampling-based methods [37, 8, 26] address this by selecting a subset of the most representative samples from the full dataset. Paul et al. [37] identify important samples very early in training. Jiang et al. [26] utilize a consistency score based on statistics collected during training. SVP [8] constructs small proxy models to find important samples. However, the diverse sample selection is not guaranteed, particularly in low data regimes, leading to suboptimal performance compared to dataset distillation (DD) methods [60]. To tackle these issues, SeiT [36] proposes a storage-efficient vision training framework by leveraging the storage efficiency of tokens. They utilize the ViT-VQGAN [56] tokenizer to extract tokens from images and store them for vision model training. As a result, they outperform other storage-efficient training methods, significantly reducing the dataset storage size.
@@
Data augmentation. Data augmentation serves as a fundamental building block to improve the model performance across diverse computer vision tasks. Various data augmentation techniques have been employed for improved performance [15, 61] or model robustness [19, 32]. Mixup [58] introduces convex combinations of pairs of images and their labels to create a smoother decision boundary. CutMix [57] replaces certain regions of an image with a patch from another image, enhancing the model’s localization ability. AutoAugment [11] leverages reinforcement learning to automate optimal augmentation policy discovery. RandAugment [12] introduces a potent policy based on random combinations of different data augmentation methods. DeiT [48] emphasizes the crucial role of data augmentation in vision transformers for improved model performance. Previous studies extensively explore the effective data augmentation policies for pixel-based vision model training. However, we observe that pixel-based data augmentations are not well-suited for tokens, posing a critical performance bottleneck. Motivated by this, we tackle the data augmentation challenge, which has not been explored in previous studies for vision model training with tokens.
@@ 3 Method
This section starts with analyzing the attributes of tokens that make pixel-based data augmentations ineffective and unsuitable, as discussed in Section 3.1. Then, we introduce our two alternative augmentation strategies, namely TokenAdapt and ColorAdapt, designed for token-based vision model training. In Section 3.2, we present the TokenAdapt module, which facilitates the application of pixel-based data augmentations to token embedding by transforming them into a compatible domain. In Section 3.3, we propose ColorAdapt, an additional color-based token augmentation, to further diversify the training data. Finally, Section 3.4 describes the overall training procedure that incorporates our augmentation strategies for vision model training.
3.1 Limited augmentations for tokens
@@
We examine why pixel-based geometric data augmentations are not well-suited for tokens, identifying two main reasons: (1) spatial information collapse during tokenization and (2) inter-token dependencies. The discrete tokenization process condenses an 2D image patch into a single -dimensional 1D vector (i.e., token embedding), causing a collapse of spatial information. This poses challenges for data augmentation, particularly for operations like horizontal flip (hFlip) that depend on the spatial details of the original pixels. We also observe that neighboring -dimensional token embeddings are interdependent, in contrast to the independence of individual pixels in an image. As a result, applying data augmentations related to interpolation (e.g., Resize or Mixup [58]) introduces undesired artifacts.
| Framework | hFlip | Mixup [58] | Affine | Top 1 Acc. |
| SeiT [36] | ✘ | ✘ | ✘ | 74.0 |
| ✔ | ✘ | ✘ | 73.2 (-0.8) | |
| ✔ | ✔ | ✘ | 72.0 (-2.0) | |
| ✔ | ✔ | ✔ | 71.3 (-2.7) |
@@
Figure 2 illustrates examples related to the failure of pixel-based augmentations for tokens. When applying a hFlip, we observe the disruption of both the diagonal line (upper row) and the cat’s silhouette (lower row) due to spatial information collapse. Specifically, only the inter-token spatial relationship is flipped, while the inner-token spatial arrangement remains unchanged. We also witness that augmentations involving interpolation can lead to undesired artifacts, such as color leaking in the case of Random Resized Crop (RRC), because of interdependence between neighboring tokens. These problems are observed in the tokens extracted from natural images as well (see Figure 3). Consequently, direct application of pixel-based augmentations to token embedding negatively affects token-based vision classifier learning as shown in Table 1.
3.2 TokenAdapt
@@
The concept of TokenAdapt is to establish a feature space that accommodates conventional pixel-based augmentations while facilitating the conversion and inversion of pre-extracted tokens under minimal computation overheads. This approach offers several advantages for our TokenAdapt. Firstly, it enables the direct reuse of well-designed pixel-based data augmentation policies. Besides, the augmented token can be easily computed via the conversion-augmentation-reverse process. Lastly, it ensures the preservation of identical feature spaces for both original and augmented tokens, allowing for the seamless utilization of existing model architectures or pre-trained weights.
We assume token inputs instead of pixels for storage efficiency [36]. Thus, the input is a sequence of tokens that represents a pixel image . Since the tokens are stored in quantized form, each token means an embedding vector in a codebook. i.e., where for a codebook with size . The goal of TokenAdapt is to design an augmentation function that converts the input tokens to augmented tokens . Note that the augmented tokens should also be quantized .
@@
Figure 4 depicts the overall pipeline of our TokenAdapt. When a token sequence and a pre-trained codebook is given, the token embedding corresponds to the codebook is obtained through lookup process, i.e., . The initial step of TokenAdapt is transforming into a representation in augmentation-compatible space . This transformation is achieved using a conversion function , i.e., . We apply augmentation at augmentation-compatible space and build an augmented representation . Then, the augmented representation is reverted to token embedding space by a reverse transformation , . At last, the augmented tokens in embedding space are quantized to codebook index as , where indicates the vector quantization process. In summary, the augmentation process of TokenAdapt is:
@@
We particularly focus on geometric data augmentations (e.g., flip, resize, affine) and Mixup [58], which result in a performance drop for the baseline method, as shown in Table 1. Figure 3 shows the difference between direct augmentation to embedding space and augmentation with TokenAdapt . We decode the augmented tokens as images to help to understand differences in embedding space. The resulting images show that TokenAdapt effectively leverages existing pixel-based augmentations on tokens, thereby mitigating the image degradation caused by their direct application.
@@
Once trained, TokenAdapt module is utilized to augment a sequence of tokens to an augmented sequence across various datasets and downstream tasks. Importantly, the augmentation-compatible feature space, not reliant on specific semantic knowledge or alignment to downstream tasks, renders it dataset-agnostic as well as task-agnostic. This inherent flexibility of our TokenAdapt module makes it highly generalizable, enabling direct application to other datasets without the need for additional training procedures.
3.3 ColorAdapt
@@
As witnessed in Cubuk et al. [11], color-based augmentations are often more effective than geometric ones in certain datasets and play an important role in data augmentation strategies. For example, color-based augmentations are particularly useful in handling diverse environmental changes caused by varying lighting (e.g., day and night) and weather conditions (e.g., rain, fog, or snow). However, integrating them with TokenAdapt poses a challenge since most color-based augmentations are originally crafted to operate exclusively in the pixel domain.
To address this, we propose ColorAdapt, a new color-based token augmentation inspired by the method [24]. Formally, our ColorAdapt function is defined as:
@@
where and represent the mean and standard deviation of each token embedding across its spatial dimension, computed for each channel.
@@
Through transforming the global statistics within each channel of token embedding, ColorAdapt mimics color-changing augmentations in a pixel domain while preserving the encoded object structure. In Figure 5, we display the images decoded from various color-augmented tokens. This illustrates how ColorAdapt influences the visual characteristics of the augmented tokens. By thoughtfully incorporating ColorAdapt into our token augmentation strategy, we improve the data diversity, exploring new opportunities to train more robust and adaptable vision models.
| Method | Input | # of images |
Top 1 Acc. |
| Full pixels | Image | 1,281 k | 109 kB | 81.8 | 140.0 GB | ||||
| Uniform random sampling | 512 k | 107 kB | 74.0 | 54.6 GB (39%) | |||||
| C-score [26] based sampling | Image | 512 k | 104 kB | 73.3 | 53.3 GB (38%) | ||||
| Adjusting JPEG quality factor to 5 | 1,281 k | 9 kB | 74.6 | 11.0 GB (8%) | |||||
| Adjusting image resolution ( 0.2) | 1,281 k | 8 kB | 75.2 | 9.6 GB (7%) | |||||
| Ours | Token | 1,281 k | 1 kB | 75.5 | 1.4 GB (1%) |
3.4 Training procedure with TokenAdapt
We incorporate the trained TokenAdapt module to enable pixel-based data augmentation to tokens. Our evaluation focuses on two crucial computer vision tasks: (1) image classification and (2) semantic segmentation. For image classification, our goal is to enhance both model performance and robustness through the incorporation of our token augmentation strategies. Specifically, we stochastically apply pixel-based data augmentation to tokens using our TokenAdapt module. For color-based augmentations, we randomly choose between Emb-Noise [36] and our ColorAdapt, with a predetermined probability. By leveraging the diverse noise from Emb-Noise and retaining reliable object structure from our ColorAdapt, we observe an overall improvement in model performance.
In semantic segmentation, we leverage our TokenAdapt for pixel-based data augmentation as its application in image classification. However, a notable distinction exists: we do not employ mixup-based augmentations in semantic segmentation due to their limited performance benefits. This behavior aligns with recent studies [55, 46, 6] that exclude mixup-based augmentation in their training recipes. Additionally, we observe that token-level Random Resized Crop (RRC) does not enhance performance in segmentation due to the importance of structural integrity in pixel-level classification tasks. Thus, we apply geometric pixel-data augmentations using only our TokenAdapt.

### v3

# SeiT++: Masked Token Modeling Improves Storage-efficient Training

@@ Abstract
Recent advancements in Deep Neural Network (DNN) models have significantly improved performance across computer vision tasks. However, achieving highly generalizable and high-performing vision models requires expansive datasets, resulting in significant storage requirements. This storage challenge is a critical bottleneck for scaling up models. A recent breakthrough by SeiT proposed the use of Vector-Quantized (VQ) feature vectors (i.e., tokens) as network inputs for vision classification. This approach achieved 90% of the performance of a model trained on full-pixel images with only 1% of the storage. While SeiT needs labeled data, its potential in scenarios beyond fully supervised learning remains largely untapped. In this paper, we extend SeiT by integrating Masked Token Modeling (MTM) for self-supervised pre-training. Recognizing that self-supervised approaches often demand more data due to the lack of labels, we introduce TokenAdapt and ColorAdapt. These methods facilitate comprehensive token-friendly data augmentation, effectively addressing the increased data requirements of self-supervised learning. We evaluate our approach across various scenarios, including storage-efficient ImageNet-1k classification, fine-grained classification, ADE-20k semantic segmentation, and robustness benchmarks. Experimental results demonstrate consistent performance improvement in diverse experiments, validating the effectiveness of our method. Code is available at https://github.com/naver-ai/tokenadapt.
@@ 1 Introduction
Recent advancements in Deep Neural Network (DNN) models [17, 50] have significantly improved their performance across various computer vision tasks. However, achieving highly generalizable and high-performing vision models demands an extensive dataset, often containing billions of data points [27, 46, 36] with large storage requirements. For example, the LAION-5B [44] dataset requires a storage size of 240TB. This storage challenge emerges as a critical bottleneck for scaling up vision models. Consequently, optimizing storage efficiency becomes important for the practical advancement of scaled-up vision models.
@@
Images require more storage than texts, due to their intrinsic complexity in converting continuous natural light signals into discrete pixels for computer representation. However, a considerable portion of image data corresponds to redundant or unnecessary details, such as repeated patterns or imperceptible details [41]. These properties lead to large storage requirements and also induce negative impacts on visual recognition tasks as mentioned in [10, 22, 35, 1, 21, 45].
Therefore, attempts have been made to reduce storage requirements by optimizing the efficiency of image data. The standard approach to reduce storage requirements is to decrease the total number of training data by removing less important samples [39, 8, 28] or condensing the data into smaller synthetic sets [62, 63]. However, as SeiT [38] illustrates, this approach can lead to a significant performance drop when compared to using the entire dataset or become impractical for large-scale datasets due to their complexity. Alternatively, some methods aim to reduce the size of each image through compression techniques or resolution adjustments [44]. However, they still face a considerable performance gap with the original dataset. Recently, SeiT [38] has made a significant breakthrough in this area by enabling the direct learning of ViTs [17] from discrete representations (i.e., tokens) pre-extracted by an offline tokenizer [59]. By using vector-quantized features instead of traditional pixel images, SeiT achieves 90% of the performance of a full storage model while requiring only 1% of the storage.
While SeiT [38] has demonstrated the potential of token-based training, exploration beyond fully supervised learning scenarios remains an untapped area. To leverage the advantages offered by large-scale data beyond human-annotated datasets, it is imperative to explore label-free learning methods. In this paper, we investigate the compatibility of SeiT with self-supervised learning approaches in token-based frameworks. Among many different self-supervised algorithms, Masked Image Modeling (MIM) [23, 58] derived from Masked Language Modeling [15] has brought significant success and has quickly become a popular approach for learning visual representations. Inspired by this, we adopt Masked Token Modeling (MTM) as a pre-training method in token-based frameworks and validate its effectiveness for storage efficiency on various benchmarks. To the best of our knowledge, this is the first approach for learning an MLM-based model directly from offline tokens, offering the potential to optimize storage efficiency for large-scale training tasks.
It is worth noting that self-supervised approaches often demand more data due to the lack of labels. MIMs also utilize data augmentation techniques to the core, including a random resized crop and horizontal flip. As previous studies [12, 11, 50] have highlighted, data augmentation plays a crucial role in large-scale model training and is a fundamental building block that is widely used in various training recipes. However, applying existing data augmentation techniques designed for pixel images to tokens poses a significant challenge due to the input domain shift. Given its importance in DNN-based models, it is crucial to address the challenges posed by applying data augmentation to tokens. This resolution holds the key to overcoming major performance bottlenecks and improving the storage efficiency of token-based frameworks.
To this end, we introduce TokenAdapt and ColorAdapt, two novel augmentation strategies designed for token-based training. TokenAdapt facilitates the integration of traditional pixel-based augmentation policies by transforming token embeddings into a feature space compatible with pixel-based augmentations. After applying augmentation to the transformed token embeddings, the augmented token embeddings are then reverted to their original token space. ColorAdapt, inspired by Adaptive Instance Normalization (AdaIN [26]), alters the color attributes of token embeddings while preserving object structure by adjusting their statistics. By integrating TokenAdapt and ColorAdapt into our token augmentation process, we enhance the effectiveness of token-based training across both fully supervised and unsupervised learning scenarios, demonstrating significant improvements in model performance.
@@
Our SeiT++ framework combines Masked Token Modeling (MTM) with the proposed token augmentation strategies, TokenAdapt and ColorAdapt. We validate the effectiveness of SeiT++ on various scenarios: (1) storage-efficient ImageNet-1k classification, (2) transfer learning on fine-grained classification, (3) ADE-20k semantic segmentation, and (4) evaluation on robustness benchmarks. As shown in Figure 1, SeiT++ consistently outperforms comparison methods in storage-efficient ImageNet-1k classification with limited storage sizes, achieving a top-1 accuracy of 77.8% with only 1.4GB of data. These results confirm that token-based training can take advantage of self-supervised learning to improve storage efficiency. We also demonstrate that our proposed token augmentations significantly contribute to storage efficiency by augmenting data effectively. We reveal our token-based augmentation methods basically yield significant boosts in supervised learning and further amplify improvements in MTM learning. Furthermore, We present how our strategies enhance the generalizability of trained models, as evidenced by our experimental results on fine-grained classification and robustness evaluation. Lastly, the applicability of our method to an alternative input format (e.g., DCT coefficients) demonstrates its extensibility.
@@ 2 Related Work
The use of discrete tokens for training vision models has recently gained attention, particularly in self-supervised representation learning. BeiT [2] employs the Masked Image Modeling (MIM) framework to recover discrete tokens from masked images and CIM [19] utilizes tokens in Corrupted Image Modeling. MAGE [32] demonstrates the incorporation of tokens from VQGAN [18] for both generative training and representation learning within a unified training framework. However, existing studies mainly focus on self-supervised representation learning and rely on online tokenization, converting images to tokens during each iteration of model training. This approach incurs notable memory consumption and computational costs, limiting the full exploitation of the storage efficiency of discrete tokens.
2.0.2 Storage-efficient Vision Training.
@@
Storage-efficient vision training aims to reduce the dataset storage size. To this end, dataset distillation (DD) [54] creates a condensed dataset by leveraging the knowledge of the original dataset. Recent research [31, 42, 43, 62, 63] highlights the effectiveness of dataset distillation in improving training efficiency. However, its high computational complexity poses challenges, particularly when applied to large-scale datasets like ImageNet-1k [14]. Sampling-based methods [39, 8, 28] address this by selecting a subset of the most representative samples from the full dataset. Paul et al. [39] identify important samples very early in training. Jiang et al. [28] utilize a consistency score based on statistics collected during training. SVP [8] constructs small proxy models to find important samples. However, the diverse sample selection is not guaranteed, particularly in low data regimes, leading to suboptimal performance compared to dataset distillation (DD) methods [63]. To tackle these issues, SeiT [38] proposes a storage-efficient vision training framework by leveraging the storage efficiency of tokens. They utilize the ViT-VQGAN [59] tokenizer to extract tokens from images and store them for vision model training. As a result, they outperform other storage-efficient training methods, significantly reducing the dataset storage size.
2.0.3 Data Augmentation.
@@
Data augmentation serves as a fundamental building block to improve the model performance across diverse computer vision tasks. Various data augmentation techniques have been employed for improved performance [16, 64] or model robustness [20, 34]. Mixup [61] introduces convex combinations of pairs of images and their labels to create a smoother decision boundary. CutMix [60] replaces certain regions of an image with a patch from another image, enhancing the model’s localization ability. AutoAugment [11] leverages reinforcement learning to automate optimal augmentation policy discovery. RandAugment [12] introduces a potent policy based on random combinations of different data augmentation methods. DeiT [50] emphasizes the crucial role of data augmentation in vision transformers for improved model performance. Previous studies extensively explore the effective data augmentation policies for pixel-based vision model training. However, we observe that pixel-based data augmentations are not well-suited for tokens, posing a critical performance bottleneck. Motivated by this, we tackle the data augmentation challenge, which has not been explored in previous studies for vision model training with tokens.
@@ 3 Method
3.1 Preliminary: Storage-efficient Vision Training (SeiT)
3.1.1 SeiT
[38] aims to develop a scalable and high-performance vision classifier with minimal storage requirements. It introduces a storage-efficient vision training framework that enables the direct learning of ViT [17] from discrete representations (i.e., tokens) pre-extracted by an offline tokenizer. In addition, SeiT further improves training by incorporating two token-specific augmentations: Token-EDA, drawing inspiration from easy data augmentation for language [55], and Emb-Noise, which introduces Gaussian noise into discrete representations. By using vector-quantized features instead of pixel images, SeiT retains 90% of the precision of a model trained on full-pixel images while using only 1% of the storage space. SeiT consists of two parts: (1) preparing the token dataset and (2) training a model using the token dataset.
3.1.2 Preparing the Token Dataset.
SeiT extracts tokens using the ImageNet-1k-trained ViT-VQGAN tokenizer [59] to prepare the token dataset for training. As illustrated in Figure 2(a), the process begins by transforming each pixel image into a sequence of tokens . Since the tokens are stored in quantized form, each token means an embedding vector in a codebook . i.e., where for a codebook with size . Here, given the tokens and the codebook , the corresponding token embeddings can be computed by lookup process. This conversion compresses each image into a more compact token sequence , where is much smaller than , thus significantly reducing storage requirements. For example, storing the full-pixel ImageNet-1k requires 140GB, but its tokenized form needs only 1.4GB. Notably, the tokens are stored on storage before model training with the codebook . Thus, during the training phase, the pre-stored tokens are used directly as input instead of the original pixel images.
3.1.3 Training Pipeline.
To train a vision classifier using tokens, SeiT initially loads the token dataset and applies Token-EDA to the tokens. The tokens are then converted to a one-hot format and randomly resized and cropped. Subsequently, the processed one-hot tokens are transformed into token embeddings using a pre-trained codebook. Techniques such as CutMix and Embedding-noise are applied to these token embeddings, and the resulting augmented tensor is fed into the model for training. Through the proposed framework, SeiT demonstrates the potential of token-based training, but it lacks exploration beyond fully supervised learning scenarios. In addition, while token-specific augmentation is proposed, the types of token augmentation are limited and the analysis of data augmentation compared to pixel-based training is insufficient.
3.2 Masked Token Modeling
In this section, we introduce Masked Token Modeling (MTM), a pre-training method in token-based frameworks without relying on labeled datasets. MTM shares a similar concept to Masked Language Modeling (MLM) [15], which works by reconstructing original tokens from their masked versions, employing an autoencoding strategy. Figure 2(b) presents an overview of how our MTM operates.
3.2.1 Masking.
We employ a variable masking ratio that follows a truncated normal distribution, as investigated in previous studies [4, 5, 32]. The tokens are randomly dropped according to the sampled masking ratio, and then the visible, unmasked token embeddings are fed into the model. By processing only a subset of token embeddings in the encoder, we significantly reduce the overall pre-training time and memory consumption, which aligns with the findings of MAE [23].
3.2.2 MTM Encoder and Decoder.
We use ViT [17] as our encoder because its architecture is capable of processing both the visible token embeddings and the original token embeddings, seamlessly accommodating their different lengths. During training, the encoder takes the visible token embeddings and outputs latent features that become the decoder input. Before decoding, the latent features are padded to the length of the original token embeddings using mask tokens M. Each mask token is a shared, learned vector indicating the presence of a missing token to be predicted. The decoder then estimates the original tokens from the padded features.
3.2.3 Training Objective.
MTM learns to accurately predict the original tokens of each masked token from the visible tokens. In particular, we use a cross-entropy loss between the decoder output and the ground truth tokens as the training objective, which is denoted by :
where denotes the predictions of the model for the masked tokens, and corresponds to the ground truth tokens for . Following MAE [23], this loss is optimized only for the masked tokens.
3.3 Data Augmentation for Tokens
While Masked Token Modeling (MTM) improves the storage efficiency of the token-based framework, there is still a performance bottleneck in improving token-based training. Since self-supervised approaches often require more data due to the lack of labels, exploring data augmentation is important to take full advantage of MTM. In this section, we analyze why existing data augmentations designed for pixel images are ineffective in the token domain. We then present TokenAdapt and ColorAdapt as solutions for token-based data augmentation.
3.3.1 Limited Data Augmentations for Tokens.
@@ 3.3 Data Augmentation for Tokens
We explore the challenges of pixel-based augmentation in the token domain and identify two main reasons: (1) spatial information collapse during tokenization and (2) inter-token dependencies. The discrete tokenization process condenses a 2D image patch into a single -dimensional 1D vector (i.e., token embedding), causing a collapse of spatial information. This collapse hinders augmentations that rely on spatial details (e.g., horizontal flip). Furthermore, unlike independent image pixels, token embeddings are inter-correlated. This complicates augmentations involving interpolation (e.g., Resize or Mixup [61]), by introducing undesired artifacts.
@@
Figure 3 shows the limitations of pixel-based augmentations in the token domain. Applying a horizontal Flip (hFlip) results in the distortion of both the diagonal line (top row) and the cat’s silhouette (bottom row) due to spatial information collapse. This occurs because only the inter-token spatial relationship is flipped, while the inner-token spatial arrangement remains unchanged. Augmentations involving interpolation can lead to undesired artifacts, such as color leakage in the case of Random Resized Crop (RRC), due to interdependence between neighboring embeddings. Therefore, we suspect that the direct application of pixel-based augmentations to token embedding may negatively affect token-based training.
3.3.2 TokenAdapt.
@@
The concept of TokenAdapt is to establish a feature space that allows for conventional pixel-based augmentations while facilitating the conversion and inversion of pre-extracted tokens with minimal computational overhead. This approach offers several advantages for our TokenAdapt. Firstly, it enables the direct reuse of well-designed pixel-based data augmentation policies. Additionally, the augmented token can be easily computed through the conversion-augmentation-reverse process. Finally, it ensures that the feature spaces for both the original and augmented tokens are preserved identically. This allows for the seamless use of existing model architectures or pre-trained weights.
@@
The goal of TokenAdapt is to design an augmentation function that converts the input tokens to augmented tokens . Note that the augmented tokens should also be quantized . Figure 4 depicts the overall pipeline of TokenAdapt. The initial step of TokenAdapt is transforming token embeddings , computed from and codebook , into a representation in augmentation-compatible space . This transformation is achieved using a conversion function , i.e., . We apply augmentation at augmentation-compatible space and build an augmented representation . Then, the augmented representation is reverted to token embedding space by a reverse transformation , . At last, the augmented tokens in embedding space are quantized to codebook index as , where indicates the vector quantization process. In summary, the augmentation process of TokenAdapt is as follows:
@@
TokenAdapt deals with geometric data augmentations (e.g., flip, resize, crop, affine) and Mixup [61]. These augmentations are commonly used in various training methods, but they can cause a performance drop in token-based training, as shown in Figure 3(b).
@@
Once trained, the TokenAdapt module is used to augment a sequence of tokens into an augmented sequence across different datasets and downstream tasks. Importantly, the augmentation-compatible feature space, which does not rely on specific semantic knowledge or alignment with downstream tasks, makes it dataset and task-agnostic. The inherent flexibility of TokenAdapt makes it highly generalizable, allowing direct application to other datasets without the need for additional training procedures.
3.3.3 ColorAdapt.
@@
As witnessed in Cubuk et al. [11], color-based augmentations are often more effective than geometric ones in certain datasets and are an important component of data augmentation strategies. For example, color-based augmentations are useful for dealing with various environmental changes caused by different lighting conditions (e.g., day and night) or weather conditions (e.g., snow). However, integrating them into TokenAdapt is challenging because most color-based augmentations are designed to work only in the pixel domain. To address this, we propose ColorAdapt, a color-based token augmentation inspired by the method [26]. Formally, the ColorAdapt function is defined as:
@@
where and are the mean and standard deviation of each token embedding across its spatial dimension, computed for each channel.
@@
ColorAdapt mimics color-changing augmentations in the pixel domain while preserving the encoded object structure by transforming the global statistics within each channel of the token embedding. Figure 5 shows the images decoded from various color-augmented tokens. This illustrates how ColorAdapt affects the visual properties of the augmented tokens. By incorporating ColorAdapt into our token augmentation strategy, we improve the data diversity, exploring new possibilities for training more robust and generalizable vision models.